# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ III.1 Kohn-Sham Equation in Plane Wave Basis
In the atomic system with periodic boundary conditions, the electronic wave functions can be expanded wit the plane waves, which take the form of
@@
where $\mathbf{G}$ and $\mathbf{k}$ respectively represent the wave vectors of plane waves and sampling points in the Brillouin zone, $n$ denotes the band index, and $\{c_{n\mathbf{k}}(\mathbf{G})\}$ are the expansion coefficients of plane wave basis. The $k$-point sampling method is typically employed as the Monkhorst-Pack scheme Monkhorst and Pack (1976). Next, the Kohn-Sham equation described with the plane wave basis can be written as
$\sum_{\mathbf{G^{\prime}}}\left[\frac{1}{2}(\mathbf{k+G})^{2}\delta_{\mathbf{GG^{\prime}}}+\tilde{V}_{\mathrm{KS}}(\mathbf{G-G^{\prime}})\right]c_{n\mathbf{k}}(\mathbf{G^{\prime}})=\varepsilon_{nk}c_{n\mathbf{k}}(\mathbf{G}),$ (32)
@@
where $f(\varepsilon_{n\mathbf{k}})$ is the Fermi-Dirac distribution function, $\rho(\mathbf{r})$ is the electron density, $V_{\mathrm{xc}}$ is the exchange-correlation potential, $E_{\mathrm{xc}}$ is the exchange-correlation energy, and $E_{\rm II}$ is the ionic energy calculated by the Ewald method Martin (2020).
According to the Hellmann-Feynman theorem Hellmann (1935); Feynman (1939), the force acting on atom $I$ is defined as Eq. 21 and the stress is defined as Eq. 23. Consequently, when including a non-local part in the norm-conserving pseudopotential, the force acting on atom $I$ of type $\tau$ can be divided into three parts, namely
@@ III.2 Iterative Diagonalization Methods
Figure 11: Preconditioned conjugate gradient algorithm for $H\psi_{i}=\lambda\psi_{i}$, where $i=1,\cdots,n$ is band index. Here $\psi_{i}$ is the NN-dimensional wave function and $H$ stands for the $N\times N$ Hamiltonian. A preconditioner $K$ is applied to the residual $H\psi_{i}-\lambda\psi_{i}$ to obtain a preconditioned gradient $g_{i}$, which is then used to construct an improved search direction $d_{i}$ by Polak-Ribière method Polak (1971). Next, the new $\psi_{i}$ and $\lambda_{i}$ are obtained Teter, Payne, and Allan (1989).
@@
In general, diagonalization of the Hamiltonian matrix is the most computationally intensive step in most KS-DFT calculations. Therefore, the development of efficient diagonalization algorithms optimized for high-performance computing (HPC) architectures is essential. In particular, recent advances in heterogeneous computing platforms have substantially accelerated scientific computations, particularly in computational chemistry and materials science Jia et al. (2013); Maintz and Wetzstein (2018); Giannozzi et al. (2020). In some cases, the plane-wave basis set is capable of dealing with systems consisting of thousands of atoms or more Gygi et al. (2006); Watanabe et al. (2017); Lucignano et al. (2019); Gonze et al. (2020); Feng et al. (2024).
ABACUS employs a unified framework that operates seamlessly across diverse hardware architectures. Key linear algebra operations, such as generalized matrix-vector multiplication (gemv) and vector division (vecdiv), are encapsulated within a platform-agnostic interface, supporting execution on platforms such as CPU, GPU, and DCU. Consequently, the algorithm achieves high computational efficiency across multiple platforms. To evaluate the efficiency of heterogeneous acceleration in ABACUS, we conducted SCF calculations using plane-wave basis sets on the Bohrium cloud platform boh (2024), comparing CPU and DCU performance at equivalent computational costs. The Davidson method was selected for these benchmarks due to its demonstrated superiority in convergence speed compared to conjugate gradient approaches. Our analysis reveals a clear scaling relationship between system size and acceleration performance, with DCU computations achieving up to 5-fold speedup for systems containing approximately 50 atoms, as illustrated in Fig. 12. This performance improvement grows increasingly pronounced in larger systems, underscoring the effectiveness of our heterogeneous computing framework.
Figure 12: Computation time per SCF step for a collection of examples (with different numbers of electrons) by using the conjugate gradient (CG) method on DCU (Red), the Davidson (DAV) method on CPU (Blue), and the CG method on CPU (Green), respectively. The line chart shows their ratio. The examples are available from the link ope (2024b). The hardware is c32_m64 for CPU calculations and 4*DCU for DCU calculations.
@@ III.3.1 Formulas
$\rho(\mathbf{r})$ $=2\mathrm{Tr}\left[\hat{f}_{H}\delta(\hat{\mathbf{r}}-\mathbf{r})\right]$ (40)
@@
The entropy is similarly formulated as:
@@
However, the use of stochastic orbitals inevitably introduces stochastic errors. To achieve higher accuracy, a large number of stochastic orbitals are typically required, which substantially increases the computational time despite its linear scaling property. For instance, fragmentation approaches have been developed to compute molecular systems Neuhauser, Baer, and Rabani (2014) and covalent materials Arnon et al. (2017); Chen et al. (2019a), while stochastic “embedding” methods have been used to calculate p-nitroaniline in water Li et al. (2019). Additionally, the energy window method proposed by Chen et al. Chen et al. (2019b) decomposes the electron density into different components based on orbital energies to reduce stochastic errors, demonstrating good generality and robustness.
For simulations of materials at extremely high temperatures, mixed stochastic-deterministic Density Functional Theory (mDFT) was proposed White and Collins (2020). The method combines the advantages of traditional KS-DFT and sDFT. The mDFT method retains the computational efficiency of sDFT for high-temperature systems while reducing stochastic errors using some of the deterministic Kohn-Sham orbitals.
@@ III.3.3 Plane-Wave-Based Implementation
Fig. 14(a) shows the parallel efficiency of sDFT and mDFT. sDFT demonstrates excellent scalability as all stochastic orbitals are independently evaluated, while mDFT involves calculating KS orbitals and additional data communication among CPU cores, reducing its efficiency compared to sDFT. Fig. 14(b) compares the operational efficiency at different temperatures by recording the average time per electronic iteration for the sDFT method. The temperature ranges from 5 to 300 eV, and 96 stochastic orbitals were adopted. In sDFT, the order of Chebyshev polynomial expansions was selected to ensure an electron error less than 1e-9. Specifically, for temperatures of 5, 10, 20, 30, 50, 100, 200, and 300 eV, Chebyshev orders of 1200, 620, 300, 220, 130, 60, 35, and 25 were used, respectively. As the temperature increases, the number of Chebyshev orders decreases, causing the wall time to drop exponentially. The sDFT and mDFT methods implemented in ABACUS has been successfully applied to study warm dense matter. Liu and Chen (2022b); Chen et al. (2024a, 2025)
@@ III.4.1 Formulas
An alternative method of KS-DFT is Orbital free density functional theory (OF-DFT) Wang and Carter (2002); Karasiev and Trickey (2012), which achieves a more affordable computational complexity of $O(N\ln N)$ or $O(N)$ by calculating the non-interacting kinetic energy directly via charge density instead of Kohn-Sham orbitals. In OF-DFT, once the kinetic energy density functional (KEDF) is defined, the total energy is a pure functional of charge density, taking the form of
@@
so that the ground state energy can be obtained by directly minimizing the total energy functional with optimization algorithms, Jiang and Yang (2004) such as the truncated Newton method Nocedal (1980) and the conjugate gradient method Hager and Zhang (2005); Gilbert and Nocedal (1992).
@@ III.4.2 Kinetic Energy Density Functional
Several analytical KEDFs have been proposed over the past few decades, Karasiev and Trickey (2012); Witt et al. (2018) and they can be categorized into two main classes. First, the local and semilocal KEDFs are characterized by their kinetic energy density as a function of the charge density, its gradient, the Laplacian of the charge density, or even higher-order derivatives thereof. Thomas (1927); Fermi (1927); Weizsäcker (1935); Luo, Karasiev, and Trickey (2018); Constantin, Fabiano, and Della Sala (2018); Kang et al. (2020) Second, the nonlocal KEDFs define the kinetic energy density at each point in real space as a functional of the nonlocal charge density. Wang and Teter (1992); Wang, Govind, and Carter (1999); Huang and Carter (2010); Mi, Genova, and Pavanello (2018); Shao, Mi, and Pavanello (2021); Sun, Li, and Chen (2023); Bhattacharjee, Jana, and Samal (2024) Semilocal KEDFs are generally more computationally efficient, while nonlocal KEDFs tend to provide greater accuracy. However, a universally applicable KEDF that effectively describes both simple metals and semiconductor systems remains largely elusive, and a systematic approach to its development remains to be explored.
@@
where $\tau_{\rm{TF}}$ is the Thomas-Fermi (TF) kinetic energy density Thomas (1927); Fermi (1927), which is accurate for free electron gas (FEG),
@@ III.4.3 Machine Learning Based Kinetic Energy Density Functional
where ${\rm{softplus}}(x)=\ln(1+e^{x})$ is an activation function commonly used in machine learning, which satisfies ${\rm{softplus}}(x)\geq 0$ and ${\rm{softplus}}(x)|_{x=\ln(e-1)}=1.$ By construction, the non-negativity constraint is satisfied:
@@
As displayed in Fig. 15, the NN for the MPN KEDF uses four descriptors four descriptors $\{\tilde{p},\tilde{p}_{\rm{nl}},\tilde{\xi},\tilde{\xi}_{\rm{nl}}\}$ as inputs. First, the semilocal descriptor $\tilde{p}$ is defined as the normalized dimensionless gradient of the charge density:
@@
where $w(\mathbf{r}-\mathbf{r}^{\prime})$ is a kernel function similar to the WT kernel function Wang and Teter (1992). This kernel function is defined in reciprocal space as:
@@
While the MPN KEDF has shown promising results for simple metals and their alloys Sun and Chen (2024a), its performance remains limited for semiconductors. To address this, Sun et al. developed a multi-channel MPN KEDF Sun and Chen (2024b) that incorporates information from multiple real-space length scales.
@@ IV.1 Numerical Atomic Orbitals
Unlike the plane-wave basis, there is no unique way to construct NAOs. For example, a popular method involves solving isolated atoms subject to confining potentials Sankey and Niklewski (1989); Porezag et al. (1995); Horsfield (1997); Junquera et al. (2001); Soler et al. (2002). Based on confining potentials, Blum et alBlum et al. (2009) proposed to generate NAOs by iteratively picking up basis functions one by one from a pool of predefined candidates, to seek the best improvement of a target energy. Alternatively, OzakiOzaki (2003); Ozaki and Kino (2004) suggested that eigenfunctions of isolated atoms in confining potentials serve as “primitive orbitals” in terms of which NAOs are expanded, and expansion coefficients can be optimized with self-consistent calculations. Apart from the above energy-based methods, another class of methods constructs NAOs towards some reference states. For example, Sanchez-Portal et al Sanchez-Portal, Artacho, and Soler (1995); Sánchez-Portal, Artacho, and Soler (1996) proposed to optimize the following “spillage” function defined as
@@
Here $S_{\mu\nu}(\mathbf{k})$ is the overlap matrix. In the original work, NAOs are chosen to be combinations of pseudo-atomic orbitals (eigenfunctions of isolated atoms with pseudo-potentials) or Slater-type orbitals. Based on the spillage formalism, Chen et al. Chen, Guo, and He (2010a) proposed to construct NAOs with localized spherical Bessel functions as basis Haynes and Payne (1997), and reference systems are chosen to be a series of isolated dimers/trimers of variable bond lengths. Recently, Lin et al. Lin, Ren, and He (2021b) further improved the basis by introducing a gradient term to the original spillage
@@
where $\hat{p}$ is the momentum operator. To use the NAO basis set in ABACUS, the collection of NAO basis sets paired with SG15 pseudopotentialsSchlipf and Gygi (2015) is publicly availableABACUS team (2020).
@@ IV.2 Kohn-Sham Equation in NAO Basis
Here, indices $i$ and $j$ run over all projects for atom $I$. Equation 82 is solved independently for each $\mathbf{k}$ point. The representation of Hamiltonian or any operator within the momentum space is obtained via a "folding" operation
@@
To solve the generalized eigenvalue problem in Eq. 82, ABACUS supports numerical libraries such as ScaLAPACK sca (2024) and the Eigenvalue SoLvers for Petaflop Applications (ELPA) elp (2024). ABACUS also provides the pole expansion and selected inversion (PEXSI) method as a low-scaling alternative (see Sec. VI.8 for details). Additionally, we have implemented GPU acceleration through specialized numerical libraries, such as the GPU-enabled ELPA implementation Yu et al. (2021). After diagonalization, the density matrix can be constructed via the formula

### v4

@@ III.1 Kohn-Sham Equation in Plane Wave Basis
In the atomic system with periodic boundary conditions, the electronic wave functions can be expanded with the plane waves, which take the form of
@@
where $\mathbf{G}$ and $\mathbf{k}$ respectively represent the wave vectors of plane waves and sampling points in the Brillouin zone, $n$ denotes the band index, and $\{c_{n\mathbf{k}}(\mathbf{G})\}$ are the expansion coefficients of plane wave basis. The $k$-point sampling method is typically employed as the Monkhorst-Pack scheme. Monkhorst and Pack (1976) Next, the Kohn-Sham equation described with the plane wave basis can be written as
$\sum_{\mathbf{G^{\prime}}}\left[\frac{1}{2}(\mathbf{k+G})^{2}\delta_{\mathbf{GG^{\prime}}}+\hat{V}_{\mathrm{KS}}(\mathbf{G-G^{\prime}})\right]c_{n\mathbf{k}}(\mathbf{G^{\prime}})=\varepsilon_{nk}c_{n\mathbf{k}}(\mathbf{G}),$ (32)
@@
where $f(\varepsilon_{n\mathbf{k}})$ is the Fermi-Dirac distribution function, $\rho(\mathbf{r})$ is the electron density, $V_{\mathrm{xc}}$ is the exchange-correlation potential, $E_{\mathrm{xc}}$ is the exchange-correlation energy, and $E_{\rm II}$ is the ionic energy calculated by the Ewald method. Martin (2020)
According to the Hellmann-Feynman theorem, Hellmann (1935); Feynman (1939) the force acting on atom $I$ is defined as Eq. 21 and the stress is defined as Eq. 23. Consequently, when including a non-local part in the norm-conserving pseudopotential, the force acting on atom $I$ of type $\tau$ can be divided into three parts, namely
@@ III.2 Iterative Diagonalization Methods
Figure 10: Preconditioned conjugate gradient algorithm for $H\psi_{i}=\lambda\psi_{i}$, where $i=1,\cdots,n$ is band index. Here $\psi_{i}$ is the NN-dimensional wave function and $H$ stands for the $N\times N$ Hamiltonian. A preconditioner $K$ is applied to the residual $H\psi_{i}-\lambda\psi_{i}$ to obtain a preconditioned gradient $g_{i}$, which is then used to construct an improved search direction $d_{i}$ by Polak-Ribière method. Polak (1971) Next, the new $\psi_{i}$ and $\lambda_{i}$ are obtained. Teter, Payne, and Allan (1989) Here, $\gamma$ controls the conjugacy of the search directions, leveraging the algorithm’s history for optimal convergence. $\theta$ acts as an optimal step size, ensuring maximal energy reduction at each iteration.
@@
In general, diagonalization of the Hamiltonian matrix is the most computationally intensive step in most KS-DFT calculations. Therefore, the development of efficient diagonalization algorithms optimized for high-performance computing (HPC) architectures is essential. In particular, recent advances in heterogeneous computing platforms have substantially accelerated scientific computations, particularly in computational chemistry and materials science. Jia et al. (2013); Maintz and Wetzstein (2018); Giannozzi et al. (2020) In some cases, the plane-wave basis set is capable of dealing with systems consisting of thousands of atoms or more. Gygi et al. (2006); Watanabe et al. (2017); Lucignano et al. (2019); Gonze et al. (2020); Feng et al. (2024)
ABACUS employs a unified framework that operates seamlessly across diverse hardware architectures. Key linear algebra operations, such as generalized matrix-vector multiplication (gemv) and vector division (vecdiv), are encapsulated within a platform-agnostic interface, supporting execution on platforms such as CPU, GPU, and DCU. Consequently, the algorithm achieves high computational efficiency across multiple platforms. To evaluate the efficiency of heterogeneous acceleration in ABACUS, we conducted SCF calculations using plane-wave basis sets on the Bohrium cloud platform, boh (2024) comparing CPU and DCU performance at equivalent computational costs. The Davidson method was selected for these benchmarks due to its demonstrated superiority in convergence speed compared to conjugate gradient approaches. Our analysis reveals a clear scaling relationship between system size and acceleration performance, with DCU computations achieving up to 5-fold speedup for systems containing approximately 50 atoms, as illustrated in Fig. 11. This performance improvement grows increasingly pronounced in larger systems, underscoring the effectiveness of our heterogeneous computing framework.
Figure 11: Computation time per SCF step for a collection of examples (with different numbers of electrons) by using the conjugate gradient (CG) method on DCU (Red), the Davidson (DAV) method on CPU (Blue), and the CG method on CPU (Green), respectively. The line chart shows their ratio. The examples are available from the link. ope (2024b) The hardware is c32_m64 for CPU calculations and 4*DCU for DCU calculations.
@@ III.3.1 Formulas
$\rho(\mathbf{r})$ $=2\mathrm{Tr}\left[\hat{f}_{H}\delta(\hat{\mathbf{r}}-\mathbf{r})\right],$ (40)
@@
The entropy is similarly formulated as
@@
However, the use of stochastic orbitals inevitably introduces stochastic errors. To achieve higher accuracy, a large number of stochastic orbitals are typically required, which substantially increases the computational time despite its linear scaling property. For instance, fragmentation approaches have been developed to compute molecular systems Neuhauser, Baer, and Rabani (2014) and covalent materials, Arnon et al. (2017); Chen et al. (2019a) while stochastic “embedding” methods have been used to calculate p-nitroaniline in water. Li et al. (2019) Additionally, the energy window method proposed by Chen et al. Chen et al. (2019b) decomposes the electron density into different components based on orbital energies to reduce stochastic errors, demonstrating good generality and robustness.
For simulations of materials at extremely high temperatures, mixed stochastic-deterministic Density Functional Theory (mDFT) was proposed. White and Collins (2020) The method combines the advantages of traditional KS-DFT and sDFT. The mDFT method retains the computational efficiency of sDFT for high-temperature systems while reducing stochastic errors using some of the deterministic Kohn-Sham orbitals.
@@ III.3.3 Plane-Wave-Based Implementation
Fig. 13(a) shows the parallel efficiency of sDFT and mDFT. sDFT demonstrates excellent scalability as all stochastic orbitals are independently evaluated, while mDFT involves calculating KS orbitals and additional data communication among CPU cores, reducing its efficiency compared to sDFT. Fig. 13(b) compares the operational efficiency at different temperatures by recording the average time per electronic iteration for the sDFT method. The temperature ranges from 5 to 300 eV, and 96 stochastic orbitals were adopted. In sDFT, the order of Chebyshev polynomial expansions was selected to ensure an electron error less than $10^{-9}$. Specifically, for temperatures of 5, 10, 20, 30, 50, 100, 200, and 300 eV, Chebyshev orders of 1200, 620, 300, 220, 130, 60, 35, and 25 were used, respectively. As the temperature increases, the number of Chebyshev orders decreases, causing the wall time to drop exponentially. The sDFT and mDFT methods implemented in ABACUS has been successfully applied to study warm dense matter. Liu and Chen (2022b); Chen et al. (2024a, 2025)
@@ III.4.1 Formulas
An alternative method of KS-DFT is Orbital free density functional theory (OF-DFT), Wang and Carter (2002); Karasiev and Trickey (2012) which achieves a more affordable computational complexity of $O(N\ln N)$ or $O(N)$ by calculating the non-interacting kinetic energy directly via charge density instead of Kohn-Sham orbitals. In OF-DFT, once the kinetic energy density functional (KEDF) is defined, the total energy is a pure functional of charge density, taking the form of
@@
so that the ground state energy can be obtained by directly minimizing the total energy functional with optimization algorithms, Jiang and Yang (2004) such as the truncated Newton method Nocedal (1980) and the conjugate gradient method. Hager and Zhang (2005); Gilbert and Nocedal (1992)
@@ III.4.2 Kinetic Energy Density Functional
Several analytical KEDFs have been proposed over the past few decades, Karasiev and Trickey (2012); Witt et al. (2018) and they can be categorized into two main classes. First, the local and semilocal KEDFs are characterized by their kinetic energy density as a function of the charge density, its gradient, the Laplacian of the charge density, or even higher-order derivatives. Thomas (1927); Fermi (1927); Weizsäcker (1935); Luo, Karasiev, and Trickey (2018); Constantin, Fabiano, and Della Sala (2018); Kang et al. (2020) Second, the nonlocal KEDFs define the kinetic energy density at each point in real space as a functional of the nonlocal charge density. Wang and Teter (1992); Wang, Govind, and Carter (1999); Huang and Carter (2010); Mi, Genova, and Pavanello (2018); Shao, Mi, and Pavanello (2021); Sun, Li, and Chen (2023); Bhattacharjee, Jana, and Samal (2024) Semilocal KEDFs are generally more computationally efficient, while nonlocal KEDFs tend to provide greater accuracy. However, a universally applicable KEDF that effectively describes both simple metals and semiconductor systems remains largely elusive, and a systematic approach to its development remains to be explored.
@@
where $\tau_{\rm{TF}}$ is the Thomas-Fermi (TF) kinetic energy density, Thomas (1927); Fermi (1927) which is accurate for free electron gas (FEG),
@@ III.4.3 Machine Learning Based Kinetic Energy Density Functional
where ${\rm{softplus}}(x)=\ln(1+e^{x})$ is an activation function commonly used in machine learning, which satisfies ${\rm{softplus}}(x)\geq 0$ and ${\rm{softplus}}(x)|_{x=\ln(e-1)}=1.$ By construction, the non-negativity constraint is satisfied
@@
As displayed in Fig. 14, the NN for the MPN KEDF uses four descriptors $\{\tilde{p},\tilde{p}_{\rm{nl}},\tilde{\xi},\tilde{\xi}_{\rm{nl}}\}$ as inputs. First, the semilocal descriptor $\tilde{p}$ is defined as the normalized dimensionless gradient of the charge density
@@
where $w(\mathbf{r}-\mathbf{r}^{\prime})$ is a kernel function similar to the WT kernel function. Wang and Teter (1992) This kernel function is defined in reciprocal space as
@@
While the MPN KEDF has shown promising results for simple metals and their alloys, Sun and Chen (2024a) its performance remains limited for semiconductors. To address this, Sun et al. developed a multi-channel MPN KEDF Sun and Chen (2024b) that incorporates information from multiple real-space length scales.
@@ IV.1 Numerical Atomic Orbitals
Unlike the plane-wave basis, there is no unique way to construct NAOs. For example, a popular method involves solving isolated atoms subject to confining potentials. Sankey and Niklewski (1989); Porezag et al. (1995); Horsfield (1997); Junquera et al. (2001); Soler et al. (2002) Based on confining potentials, Blum et al. Blum et al. (2009) proposed to generate NAOs by iteratively picking up basis functions one by one from a pool of predefined candidates, to seek the best improvement of a target energy. Alternatively, Ozaki Ozaki (2003); Ozaki and Kino (2004) suggested that eigenfunctions of isolated atoms in confining potentials serve as “primitive orbitals” in terms of which NAOs are expanded, and expansion coefficients can be optimized with self-consistent calculations. Apart from the above energy-based methods, another class of methods constructs NAOs towards some reference states. For example, Sanchez-Portal et al. Sanchez-Portal, Artacho, and Soler (1995); Sánchez-Portal, Artacho, and Soler (1996) proposed to optimize the following “spillage” function defined as
@@
Here $S_{\mu\nu}(\mathbf{k})$ is the overlap matrix. In the original work, NAOs are chosen to be combinations of pseudo-atomic orbitals (eigenfunctions of isolated atoms with pseudo-potentials) or Slater-type orbitals. Based on the spillage formalism, Chen et al. Chen, Guo, and He (2010a) proposed to construct NAOs with localized spherical Bessel functions as basis, Haynes and Payne (1997) and reference systems are chosen to be a series of isolated dimers/trimers of variable bond lengths. Recently, Lin et al. Lin, Ren, and He (2021b) further improved the basis by introducing a gradient term to the original spillage
@@
where $\hat{p}$ is the momentum operator. To use the NAO basis set in ABACUS, the collection of NAO basis sets paired with SG15 pseudopotentials Schlipf and Gygi (2015) is publicly available. ABACUS team (2020)
@@ IV.2 Kohn-Sham Equation in NAO Basis
Here, indices $i$ and $j$ run over all projects for atom $I$. Equation 82 is solved independently for each $\mathbf{k}$ point.
In the context of LCAO-based DFT calculations within the ABACUS package, real-space uniform grid integrals hold significant importance as they are key to multiple critical computational steps. For example, the grid integration is used to construct the local potential term of the Hamiltonian and the electron density in real space, as well as the Pulay force term when calculating the atomic forces. To optimize these grid integrals, we partition real-space grids across MPI processes and use parallel matrix operations, which substantially accelerate the overall computation. Zhang et al. (2024b)
The representation of Hamiltonian or any operator within the momentum space is obtained via a "folding" operation
@@
To solve the generalized eigenvalue problem in Eq. 82, ABACUS supports numerical libraries such as ScaLAPACK sca (2024) and the Eigenvalue SoLvers for Petaflop Applications (ELPA). elp (2024) ABACUS also provides the pole expansion and selected inversion (PEXSI) method as a low-scaling alternative (see Sec. VI.9 for details). Additionally, we have implemented GPU acceleration through specialized numerical libraries, such as the GPU-enabled ELPA implementation. Yu et al. (2021) After diagonalization, the density matrix can be constructed via the formula